# GATSY: Graph Attention Network for Music Artist Similarity thanks:

### v2

@@ V-D Artist Recommendation
To answer RQ5, we now describe how our method and in general the artist similarity task can work as a recommender system. To enhance heterogenous recommendations in terms of music genres, in this Section we show some of the results we had from the unsupervised experiments. These results provide also a qualitative perspective on how these systems work in practice.
@@
Recommendation for fictitious artists. One of the novelties of our work relies on the possibility of adding fictitious artists like in [44] to create a bridge between two possibly distant genres and still obtain plausible suggestions. We trust this procedure thanks to the robustness of GATSY to the input features.
Let us assume that we want to augment the dataset with an artist that is not real. We need to provide it with a feature vector $x_{fic}$ and add a row to the adjacency matrix. The new row is a binary vector $a_{i}$ defined as follows:
@@
To illustrate the performances with the non-existing artists, we have conducted experiments with different mixings in $\mathcal{S}$. In particular, they were considered combinations among the same or similar music genres, referred to as Likely mixings (Table VI), and very different ones Unlikely mixings (Table VII). As expected, the likely mixings keep a consistency, in terms of music genres, all over the recommendations, which is not the case for the unlikely ones. Interestingly in the latter scenario, the recommendations converge on “Dance/Jazz” artists. The last artist Arthur Baker is also a famous hip-hop/rap producer, which corroborates the validity of our recommendation.
@@ VI CONCLUSION
In this work, we addressed the task of music artist similarity by proposing GATSY, a graph neural network framework that is robust to the absence of input features, and more expressive and efficient than previous approaches. We developed an augmented version of the Olga dataset that now includes labels for the artists, allowing us to enhance the unsupervised dataset and implement a genre-aware model. However, we provide evidence that using genre-based supervision does not improve the artist similarity performance, as two artists can be similar even if they come from different genres and also because is hard to identify a unique genre for a music artist. We also showed in Section V-D an artist recommendation pipeline, that fits more within real-world scenarios. We introduced the possibility of creating “bridge” nodes to connect unliked artists. By performing this data augmentation via node injection, it is possible to connect two or more previously unrelated artists and generate recommendations based on the new topology. Many challenges still need to be solved by the research community. In future works, it would be interesting to assess the capabilities of GATSY in long-tail recommendation, and evaluate whether using random features when hand-crafted are not available could be part of the solution. Moreover, other challenges like the absence of benchmark datasets should be of interest to the community. Even the development of a more effective and scalable model, that can satisfy the demand coming from modern music streaming services.

### v3

@@ V-D Artist Recommendation
To answer RQ5, we now describe how our method and, in general, the artist similarity task can work as a recommender system. To enhance heterogenous recommendations in terms of music genres, in this Section we show some of the results we had from the unsupervised experiments. These results also provide a qualitative perspective on how these systems work in practice.
@@
Recommendation for fictitious artists. One of the novelties of our work relies on the possibility of adding fictitious artists like in [42], to create a bridge between two possibly distant genres and still obtain plausible suggestions. We trust this procedure thanks to the robustness of GATSY to the input features.
Let us assume that we want to augment the dataset with an artist who is not real. We need to provide it with a feature vector $x_{fic}$ and add a row to the adjacency matrix. The new row is a binary vector $a_{i}$ defined as follows:
@@
To illustrate the performances with the non-existing artists, we have conducted experiments with different mixings in $\mathcal{S}$. In particular, they were considered combinations among the same or similar music genres, referred to as Likely mixings (Table VI), and very different ones Unlikely mixings (Table VII). As expected, the likely mixings keep a consistency, in terms of music genres, all over the recommendations, which is not the case for the unlikely ones. Interestingly in the latter scenario, the recommendations converge on “Dance/Jazz” artists. The last artist, Arthur Baker, is also a famous hip-hop/rap producer, which corroborates the validity of our recommendation.
@@ VI CONCLUSION
In this work, we addressed the task of music artist similarity by proposing GATSY, a graph neural network framework that is robust to the absence of input features, and more expressive and efficient than previous approaches. We developed an augmented version of the Olga dataset that now includes labels for the artists, allowing us to enhance the unsupervised dataset and implement a genre-aware model. However, we provide evidence that using genre-based supervision does not improve the artist similarity performance, as two artists can be similar even if they come from different genres, and also because it is hard to identify a unique genre for a music artist. We also showed in Section V-D an artist recommendation pipeline that fits more within real-world scenarios. We introduced the possibility of creating “bridge” nodes to connect unliked artists. By performing this data augmentation via node injection, it is possible to connect two or more previously unrelated artists and generate recommendations based on the new topology. Many challenges still need to be solved by the research community. In future works, it would be interesting to assess the capabilities of GATSY in long-tail recommendation and evaluate whether using random features when hand-crafted features are not available could be part of the solution. Moreover, other challenges, like the absence of benchmark datasets, should be of interest to the community. Even the development of a more effective and scalable model that can satisfy the demand coming from modern music streaming services.